OpenStudy (kainui):

Why is ((y-2x)/(2y-x)) equal to ((2x-y)/(x-2y))?

OpenStudy (anonymous):

\[(y-2x)\div(2y-x)=(2x-y)\div(x-2y)\] cross multiply it....this is what you get \[(y-2x)(x-2y)=(2x-y)(2y-x)\] and then solve it

OpenStudy (king):

just ((y-2x)/(2y-x)) multiply by minus

OpenStudy (king):

sry just mutiply ((y-2x)/(2y-x)) by minus

OpenStudy (anonymous):

xD 3 years ago

OpenStudy (anonymous):

simply ((y-2x)/(2y-x)) = (-(2x-y)/-(x-2y)) = (2x-y)/(x-2y)

OpenStudy (anonymous):

got it :D

OpenStudy (kainui):

Hey get out of here! This is my past we're talking about hahaha